A 0.75kg mass moving with a velocity of v,=1.2 –3.4j m/s , strikes a 2.4kg mass moving with a velocity of v2=0.58 jm/s . Assua. 0.466 m/s; b. – 52.1°
Answer #1

a) let the speed of the combined mass after the collision is v

Using conservation of momentum

0.75 * (1.2 i - 3.4 j) + 2.4 * (0.58 j) = (0.75 + 2.4) * v

solving for v

v = 0.29 i - 0.37 j

speed of combined mass = sqrt(0.29^2 + 0.37^2)

speed of combined mass = 0.466 m/s

the speed of combined mass is 0.466 m/s

b)

for the angle

angle of the trajectory = atan(-0.37/0.29)

angle of the trajectory = - 52.1 degree

the angle of trajectory is - 52.1 degree